# DoubleAgents: Interactive Simulations for Alignment in Agentic AI

### v2

@@ 4.1.2. Results and Findings
As summarized in Table 1, our summary-based context approach achieves the highest F1 score for predicting relevant policies given a specific problem state. The improvement is statistically significant when compared to all three baseline conditions, with an average F1 score of 0.70, demonstrating the effectiveness of using structured progress summaries to guide policy selection. Notably, simply aggregating speaker- and slot-related context without summarization does not improve performance; the combined-context baseline yields only 0.27 average F1, indicating that raw context alone is insufficient for accurate policy retrieval. Based on these results, we adopt the summary-based approach as the default mechanism for selecting policies, which are then passed to the orchestrator agent to guide plan generation in a policy-aligned manner.
Prompting Techniques F1 Score
Mean SD
@@
Slot-only .11 .038
Slot + Speaker .27 .017
@@
Table 1. The technical evaluation results for Policy Section comparing four prompting techniques. The best scores are shown in bold. Statistically significant improvements (independent two-sample t-test, $p<0.001$) over all baselines are marked with ***.
Prompting Techniques Accuracy
Mean SD
@@
Few-shot w Policy .93** .015
Table 2. The technical evaluation results for Edge Case Detection comparing different prompting techniques. The best scores are shown in bold. Statistically significant improvements (independent two-sample t-test, p ¡ 0.01) over both baselines are marked with **, improvements over one baseline are marked with †.
4.2. Edge Case Detection Experiments
4.2.1. Experiment Settings
Edge case detection involves identifying whether a respondent’s message presents a scenario that falls outside the coverage of existing coordination policies. For example, a speaker requesting to present via Zoom, when no policy has addressed virtual presentations, should be flagged to the user for clarification and further instruction. To enable reliable detection of such edge cases, we adopt a few-shot prompting strategy with access to existing policies (denoted as ”few-shot w/ policy”). We compare this approach against two baselines: (1) zero-shot prompting without any demonstration examples, and (2) few-shot prompting without access to policy references (denoted as ”few-shot w/o policy”). Concrete prompt templates for each setup are provided in Appendix D.2. To evaluate performance, we constructed a test suite of 100 examples, manually validated by the lead authors to ensure coverage and quality. The construction methodology for this test suite is detailed in Appendix D.2. All experiments were conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100. We report the average accuracy across three independent runs for each condition.
4.2.2. Results and Findings
As shown in Table 2, our few-shot prompting with policy method significantly outperforms both baselines, with statistically significant improvements across evaluation runs. These results demonstrate the effectiveness of incorporating policy context into the detection process, enabling the model to first analyze existing policies before determining whether a given case falls outside their scope. Together, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection. This ensures that novel or ambiguous cases are reliably flagged to users, enabling timely clarification and maintaining a high-quality user experience.
@@ 5.1. Goal and Overview
To investigate how users engaged with DoubleAgents’ features and how interactive simulation shaped their use of agentic behavior, we recruited 10 participants to use DoubleAgents across two separate sessions. Our goal was to assess users’ perceived system usefulness, their understanding and reliance on the system, and how their understanding and calibration of agent behavior evolved over time. During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to participants’ values and preferences via their policies, email templates, and stop hooks. Participants completed two questionnaires. First, an AI comfortableness survey measured trust and alignment needs for proactive actions and comfort with automating planning and execution, administered before Session 1, after Session 1, and after Session 2. Second, a System perception survey captured UX and perceptions of 15 system features after Session 1 and after Session 2, enabling us to track how repeated use shifted alignment strategies, delegation boundaries, and reliance. Detailed study procedures and data analysis methods are provided in Appendix A.
5.2. Participants
@@
We recruited 10 participants (7 female, 3 male; aged 21–30, M = 25.7) from our university community. All participants (10 of 10) reported using generative AI automation tools daily. Participants primarily used tools such as AI writing assistants (10 of 10), AI coding assistants (7 of 10), and AI design/creative tools (5 of 10). All participants reported solid understanding of generative AI: 7 described having in-depth knowledge of models and tools, while 3 reported a good general understanding of key concepts. Participants’ experience with human coordination tasks varied widely. 2 participants had moderate to extensive experience, having led team schedules or group planning multiple times. 3 participants had some experience, having organized small tasks or events independently. 4 participants had minimal experience, primarily helping or shadowing others in coordination tasks. Finally, 1 participant had no prior experience, having never hosted or coordinated any tasks.
5.3. Result
5.3.1. Consistently Positive User Experience
Across both sessions, all 10 participants reported a consistently positive experience with DoubleAgents. Overall system ratings remained high and showed a slight improvement (from 6.30/7 to 6.62/7). Participants also perceived the system as effective in generating optimal plans, actions, and emails (5.90/7 to 5.88/7), and judged it as robust (5.60/7 to 5.88/7). Both mental and physical demand remained low across sessions, decreasing from 2.90/7 to 2.13/7 and from 1.50/7 to 1.13/7, respectively. Participants reported high perceived performance (6.30/7 to 6.75/7), low frustration (1.80/7 to 1.25/7), and a near-significant reduction in effort (3.20/7 to 1.88/7, $p=0.054^{{}^{\wedge}}$). Together, these results indicate a consistently positive and efficient user experience across sessions, with effort showing the strongest improvement.
5.3.2. Increasing User Reliance Driven by DoubleAgents’ Design
Despite largely stable system-level scores, participants’ perceived reliance on DoubleAgents increased significantly over sessions (from 4.90/7 to 5.75/7, $p=0.02^{*}$). Participants characterized this shift not as “unconditional reliance”, but as increased comfort with delegation as the system’s behavior became more user-aligned, legible, and shapeable—enabled by clearer controls, predictable execution patterns, and explicit points of intervention. A correlation analysis between reliance and feature-level perceptions (Spearman; FDR-corrected) showed that reliance was associated with features supporting control and explainability, including state visualization ($r=0.78$, $p<0.05^{*}$) and confirmation-based user control ($r=0.51$, $p<0.05^{*}$). For example, P6 found the Communication Visualization section of the interface useful in showing the full email exchange histories between the organizer and each speaker. She mentioned, “Everything is consistent inside the system—when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication Visualization on the side.” Increases in reliance over time were strongly linked to DoubleAgents’ human-AI alignment features. In particular, reliance increased with perceived system robustness demonstrated through interactive simulation ($r=0.54$, $p<0.05^{*}$), and was reinforced by repeated successful issue flagging via stop hooks ($r=0.65$, $p<0.01^{**}$) and perceived value alignment through coordination policies and email writing ($r=0.54$, $p<0.05^{*}$). Reliance was negatively correlated with mental demand (r=0.47r=-0.47, $p<0.05^{*}$) and showed no significant correlation with other system experience or perception items. Collectively, these findings suggest that reliance increased not merely due to perceived system capability or general UX, but because of DoubleAgents’ support for alignment in design. The following section details how participants experienced and leveraged this alignment through interactive simulation.
Figure 7. Average user perception scores toward comfort with automating the (a) planning and (b) execution stage in seminar organizing, across three stages (average scores across 10 users: before Session 1, after Session 1, and after Session 2). denotes statistical significance at $p<0.1$, * at $p<0.05$, ** at $p<0.01$, *** at $p<0.005$.
5.3.3. Increasing User Comfort with AI Automating Human Coordination Tasks: Contributing Alignment Supports.
Over interactive simulations, participants became more comfortable delegating both planning and execution tasks to agentic AI. For planning (Figures 7a, participants became more comfortable letting the AI visualize current status (4.80/7 to 6.50/7, $p<0.05^{*}$), propose time assignments (5.10/7 to 6.38/7, $p<0.1^{\vee}$), decide daily next steps (4.70/7 to 6.50/7, $p<0.005^{***}$), and monitor/flag issues (5.50/7 to 6.50/7, $p<0.01^{**}$). For execution, comfort increased for drafting-only emails across all four types (from around 6.0 before to nearly 7.0/7 after, $p<0.01^{\vee}$), and also increased for draft-and-send tasks on users’ behalf—e.g., initial outreach rose from around 2.7/7 to 4.6/7 ($p<0.05^{*}$), with similar gains for follow-ups and final confirmation ($p<0.1^{\vee}$).
Participants attributed this increased comfort to DoubleAgents’ alignment supports: interactive simulation; user-editable coordination policies that guide next-step decisions; email templates and style controls that realize those actions; and stop hooks for issue escalation. Together, these supports advanced bidirectional alignment: the AI agent’s actions became better aligned with users’ needs and values, and users developed greater understanding and trust, becoming more comfortable with the AI taking initiative in automating this human coordination task.
Interactive simulation as a testbed for policy, email, and stop-hook calibration. Simulations supported alignment by letting participants safely probe edge cases, identify where the agent’s behavior diverged from expectations, and iteratively refine (1) what the agent should do (coordination policies), (2) how it should carry out those actions in email (templates and style), and (3) when it should pause for confirmation (stop hooks). Participants emphasized that interactive simulations increased confidence in their perception of the system’s robustness (5.80/7 to 6.38/7). They found simulated test cases realistic and representative of real-life scenarios (5.90/7 to 6.75/7), and especially useful for surfacing potential problems (5.80/7 to 6.50/7). P3 described how simulation exposed realistic exceptions and prompted advance policy-setting: “The simulations [over the 2 days] felt realistic and true to the person [they] were simulating, like their personality and emails… The simulation elicited possible problems that would happen in real life… Like [having] people asking for funding or [needing] to reschedule—it forced me to think about how I would handle that in real life and to create relevant policies, and that was good. I can trust the system to handle that in the future for me if scenarios like that occur, since it already happens in the simulation.
Email templates and coordination policies fostered increased comfort with AI automation. Participants’ comfort with AI automating planning and execution increased over time as the system’s actions became more consistent, better grounded in the users’ own values and rationale, and increasingly “in their voice.” Carrying over value-aligned coordination policies made participants more comfortable delegating planning—e.g., having the AI decide daily next steps, track and visualize progress, make assignment decisions, and monitor/escalate issues when cases did not fit existing policies (Figure 7a). Email-template carryover particularly supported execution delegation. Comfort increased for drafting-only emails and for draft-and-send tasks (Figure 7b). One exception was drafting and sending clarification emails without any user confirmation, which showed little change (3.00/7 to 3.90/7 to 3.90/7). P10 noted that clarifications often require nuanced, user-specific judgment: “For clarification emails, I might want to double-check… [they’re] usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI…” As a result, participants preferred to remain in the loop for clarification messages. Overall, participants described an early sense of being understood and represented by the system, as if it “got them,” when iterating on email templates and coordination policies. For example, P4 noted that edits to greetings, closings, and phrasing carried over into later drafts: “Compared to the last [usage], DoubleAgents get my style… so I edited fewer times this time.” Similarly, after refining policies, P3 observed that the agent’s choices matched their expectations: “I found that the agent’s recent actions followed exactly the policies I had in mind—super aligned.” This growing alignment helped the system reflect their intended actions and communication style, increasing their comfort with offloading tasks to AI agents.
@@ 5.3.2. Increasing User Reliance Driven by DoubleAgents’ Design
Stop hooks helped establish safe delegation boundaries, fostering users’ comfort with delegation. Participants highlighted stop-hook flagging as particularly important for maintaining safe delegation boundaries. For monitoring and surfacing potential issues, P1 described issue flagging as “one of the most important features of the system” and shared, “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things—and [then DoubleAgents] respects that…” P8 described a case where flagging prevented a misinterpretation of progress—when they skimmed the respondent agent’s reply and thought everything was resolved, the flag showed that it was not. Thus, with these “stop hooks” continuously monitoring for exceptions and prompting intervention when needed, participants felt much more comfortable offloading tasks to the AI.
Figure 8. Average user perceptions of alignment needs, expectations, and trust in agentic AI systems (average scores across 10 users: before Session 1, after Session 1, and after Session 2). * denotes statistical significance at $p<0.05$ and ** at $p<0.01$.
Increasing Expectations and Needs for Alignment in Agentic AI
Across sessions, participants reported stronger expectations for how proactive AI in general should behave and align with them (Figure 8). In the AI Comfortableness Survey, participants reported a significant increase in wanting the AI’s actions to align with their preferences, values, and priorities, from 6.10/7 (before using the system, based on their previous AI experiences) to 7.00/7 after two uses of DoubleAgents ($p<0.05^{*}$). Participants’ desire to stay in control remained high (5.30 to 5.75). They also reported an increased need for detailed system rationales and explanations before the system acts, from 5.00/7 to 6.00/7 ($p<0.05^{*}$). Participants framed clear and abundant explanations as a core alignment mechanism for building an accurate mental model of the AI’s decision process, enabling verification, and determining when and how to intervene. P6 mentioned, “[With the system explanation] I felt like I could tell exactly what it’s doing at every stage so there’s a lot of transparency… if I didn’t get that, I would feel lost [and] not know what’s going on…” We interpret this as evidence that DoubleAgents’ interaction design helped participants recognize the importance of alignment and articulate stronger alignment expectations for future proactive AI usage.
Furthermore, participants also reported statistically significant increases in trust to let similar AI systems act on their behalf, from 3.20/7 to 5.00/7 ($p<0.01^{**}$), even as their assessments of AI reliability remained relatively the same. This suggests that alignment features can increase comfort with delegation beyond improvements in perceived raw capability. P1 noted that even though DoubleAgents represented only one specific experience, “I felt more optimistic… [and] in seeing similar systems like DoubleAgents in the future, I would generally trust such AI systems more.”.

### v3

@@ 4.1.2. Results and findings
As shown in Table 1, the summary-based context yields the highest F1 score for predicting relevant policies given a specific problem state, outperforming all other input conditions. The improvement is statistically significant compared to the three baselines, indicating that providing a structured, high-level representation of task progress leads to more accurate policy selection. In contrast, simply exposing the model to raw contextual information, either from speaker personas, slot availability, or their combination, results in substantially lower performance. Notably, the combined-context condition does not improve over simpler variants, suggesting that aggregating low-level context without abstraction is insufficient for effective policy retrieval. Based on these findings, we adopt the summary-based representation as the default input for policy selection in DoubleAgents. The selected policies are then passed to the coordination agent to guide plan generation in a policy-aligned manner.
4.2. Edge Case Detection Experiments
4.2.1. Experiment settings
Edge case detection aims to determine whether a respondent’s message introduces a scenario that falls outside the scope of existing coordination policies. For instance, if a speaker requests to present via Zoom while no policy addresses virtual presentations, the system should flag this case for user clarification and further instruction.
To study how different prompting strategies affect detection performance, we compare three conditions that vary in the level of guidance and policy awareness provided to the model: (1) zero-shot prompting, where the model is given only the task instruction and the incoming message without any examples; (2) few-shot prompting without policy context (few-shot w/o policy), where the model is provided with representative examples of edge and non-edge cases but does not have access to the current set of coordination policies; and (3) few-shot prompting with policy context (few-shot w/ policy), where the model is given both demonstration examples and the existing policies, allowing it to assess whether a new situation is already covered or requires escalation.
Prompt templates for all conditions are provided in Appendix E.2. We construct a test suite of 100 examples, manually validated by the lead authors to ensure diversity and coverage across common coordination scenarios and edge cases. The construction methodology is detailed in Appendix E.2. All experiments are conducted using GPT-4o, with temperature set to 0.7 and a maximum token limit of 100. We report the average accuracy across three independent runs for each condition.
4.2.2. Results and findings
As shown in Table 1, the few-shot prompting with policy context achieves the highest accuracy among the three conditions, outperforming both zero-shot and few-shot w/o policy baselines with statistically significant improvements. These results suggest that access to explicit policy context enables the model to more reliably determine whether a given scenario is already covered or requires escalation.
In comparison, zero-shot prompting struggles to consistently distinguish between routine variations and true edge cases, while few-shot prompting without policy context benefits from examples but lacks the ability to ground its decisions in the current policy set. This often leads to either over-flagging benign cases or missing subtle policy gaps.
Based on these findings, DoubleAgents adopts the policy-aware few-shot approach as its default mechanism for edge case detection. By grounding decisions in both prior examples and explicit policy coverage, the system more reliably surfaces novel or ambiguous situations, enabling timely user intervention while maintaining a smooth coordination workflow.
Prompting Techni. Mean SD
@@ 4.2.2. Results and findings
TimeSlot-only .11 .038
TimeSlot + Speaker .27 .017
@@
(a) Policy Selection (F1 Score).
Prompting Techni. Mean SD
@@
Few-shot w/ Policy .93** .015
(b) Edge Case Detection (Accuracy)
Table 1. Technical evaluation results. Best scores in bold. Significance via independent two-sample t-test: † $p<0.05$ over one baseline, ∗∗ $p<0.01$, ∗∗∗ $p<0.001$ over all baselines.
@@ 5.1.1. Overview
To investigate how users and AI can reach alignment through repeated interaction and refinement — and how this alignment enables comfortable offloading — we recruited 10 participants across two sessions over two consecutive days. Each participant completed approximately 2.5 hours total: a 30-minute preparation and setup phase, followed by a 60-minute usage and evaluation session each day. Our goal was to assess whether comfort with delegation evolves over time, how the system’s three distributed cognition components support users in learning what is safe to delegate, and where users still need to maintain control. During each session, participants worked in a simulated environment to schedule a seminar with four dates and four speakers, while the system gradually adapted to participants’ values and preferences through their policies, email templates, and stop hooks. Participants completed two questionnaires. An AI perception survey measured alignment expectations, trust, and comfort with automating planning and execution; it was administered before Session 1, after Session 1, and after Session 2. A DoubleAgents experience survey captured UX, task load, and perceptions of system features after each session. Together, these instruments allowed us to track how repeated use shifted alignment, delegation boundaries, and the evolving balance between offloading and control. Detailed study procedures and data analysis methods are provided in Appendix C.
Figure 3. Average user-reported scores across 10 participants for comfort with automating (a) planning and (b) execution tasks, across sessions over time (1: strongly disagree, 7: strongly agree).  $p<0.1$,  $p<0.05$, ∗∗ $p<0.01$, ∗∗∗ $p<0.005$.
5.1.2. Participants
@@
We recruited 10 participants (7 female, 3 male; aged 21–30, $M=25.7$) from our university community via personal contacts and departmental mailing lists. All reported daily use of generative AI tools, including AI writing assistants (10/10), coding assistants (7/10), and design tools (5/10). Seven described in-depth knowledge of generative AI models, while three reported a good general understanding of key concepts. Participants’ experience with human coordination tasks varied widely: 2 had moderate to extensive experience, having led team schedules or group planning multiple times; 3 had some experience, having organized small tasks or events independently; 4 had minimal experience, primarily helping or shadowing others in coordination tasks; and 1 had no prior experience.
5.2. Results
We organize our findings around the central question of how humans and AI reach alignment through repeated interaction and refinement, enabling comfortable offloading. We examine three aspects: whether comfort with delegation evolves over time (§5.2.1), how the three distributed cognition components support participants in learning what is safe to delegate (§5.2.2), and where participants still need to maintain control (§5.2.3).
5.2.1. Did participants become more comfortable delegating control over time?
Participants reported increased comfort delegating both planning and execution tasks to AI
For planning (Figure 3a), comfort ratings rose across all four sub-tasks (letting the AI track and visualize task state, solve scheduling constraints such as time-slot assignments, reason about context to determine next steps, and monitor for edge cases requiring user attention), with averages climbing from roughly 5/7 before deployment to around 6.5/7 afterward. Constraint solving for slot allocation is a comparatively task-specific capability and showed only marginally significant improvement. In comparison, the planning sub-tasks, tracking, monitoring, and flagging shows strong comfort gains over time, consistent with the distributed-cognition affordances of the coordination agent that help organizers offload ongoing awareness work.
For execution (Figure 3b), the pattern split by autonomy level: participants were already fairly comfortable letting the system draft emails on their behalf, and comfort rose further (from around 5.7/7 to nearly 6.8/7 across all four email types, each at least marginally significant). By contrast, comfort with the system both drafting and sending emails—where the social stakes of an error are higher—started considerably lower and remained so, though it still improved meaningfully (from roughly 3/7 to 4.3/7 across email types). Among the four email types, initial outreach showed strong gains, likely because its content is relatively formulaic and low-variance; follow-up and confirmation emails improved modestly, while clarification emails showed no significant change over the alignment process—an exception we return to in Section 5.3.2.
Participants’ perceived reliance on DoubleAgents, as measured by the DoubleAgents Experience survey, increased significantly (4.90/7 to 5.75/7, $p=0.02^{*}$). Participants characterized this not as unconditional reliance, but as increased comfort with delegation as the system’s behavior became more predictable and legible — feeling both aligned with how they wanted to work and comfortable letting the system act on their behalf. P2 articulated this arc explicitly: “I would intervene more at first but if it can demonstrate that it’s consistently able to do it correctly, I’m going to let it act on its own on my behalf more autonomously.”
5.2.2. How did participants use and rate the three distributed cognition components?
Policies and templates helped participants encode their values, reducing effort over time.
Carrying over user-defined coordination policies strengthened alignment in planning: participants found the AI’s decisions on next steps, progress tracking, and escalation increasingly reflective of their intent. Email-template carryover similarly improved alignment in execution, with participants reporting that drafts better matched their expectations over successive sessions. Quantitative measures supported this pattern: perceived value alignment via policies and templates increased from 5.50/7 to 6.12/7, drafting comfort rose from approximately 6.0 to nearly 7.0/7 across all email types, and effort dropped from 3.20/7 to 1.88/7 ($p=0.054^{\wedge}$). P4 noted that edits to greetings, closings, and phrasing carried over into later drafts: “Compared to the last [usage], DoubleAgents get my style… so I edited fewer times this time.” P3 observed: “I found that the agent’s recent actions followed exactly the policies I had in mind — super aligned.
Visualizations made the agent’s reasoning legible, enabling confident evaluation.
The dashboard’s temporal, social, and procedural views let participants verify what the agent was doing without parsing raw data. P6 found the communication visualization useful for showing full email exchange histories: “Everything is consistent inside the system — when some speakers were not followed up, I could map it back to the visualization and confirm accuracy… [It] helped me build trust… For tomorrow, I’ll spend less time checking the email replies [inside the chat interface] one by one, and instead [just check] the Communication History on the side.
The coordination agent served as memory and reasoning partner, removing tracking burden.
By maintaining full state across days and weeks of coordination—who was contacted, who responded, what is pending—the agent freed participants from the cognitive overhead of tracking progress themselves. This let them focus on judgment and decision-making rather than bookkeeping. P7 noted that DoubleAgents was “pretty good at figuring out plans, keeping track of who’s responded, who’s not, who’s available, who’s not, like what to do moving forward.” P1 similarly shared that “now …I didn’t have to use my brain too much [to make sure things are consistent], just a little bit—so it handled the task really well.
5.2.3. Where did participants still need control rather than full delegation?
While comfort with delegation grew, participants identified boundaries where human judgment remained essential.
Stop hooks for flagging edge cases.
@@ 5.2.3. Where did participants still need control rather than full delegation?
Participants highlighted stop-hook flagging as critical for maintaining safe delegation boundaries. P1 described issue flagging as “one of the most important features of the system”: “[With the issue flagging,] I skim [through the other features] now… It made me trust the system because I know it won’t just execute something random… It flags when it doesn’t know… That helps elicit my values and how I want to handle things — and [then DoubleAgents] respects that…” P8 described a case where flagging prevented a misinterpretation of progress: when they skimmed a reply and thought everything was resolved, the flag showed that it was not. Stop hooks gave participants confidence to offload routine execution precisely because they knew the system would pause at points of uncertainty. Stop hooks gave participants confidence to offload routine execution precisely because they knew the system would pause at points of uncertainty. Comfort in offloading, in other words, was built not by the agent performing flawlessly, but by making its boundaries of confident action transparent and actionable.
Clarification emails as a natural boundary for delegation.
The notable exception to increasing delegation comfort was clarification emails sent without user confirmation, which showed little change across sessions ($p>0.1$). P10 explained: “For clarification emails, I might want to double-check… [they’re] usually not in the plan [or policies]… I’m not comfortable [offloading that] to AI…” Unlike routine emails that follow established templates, clarification emails require the agent to identify ambiguity, judge what information is missing, and formulate the right question — all of which demand user-specific judgment that resists encoding into policies. This suggests a natural boundary for delegation: as agents take on more execution, the critical human role shifts toward decision-making at points of uncertainty, precisely the moments where knowing “when and how to ask the user” becomes the key alignment challenge.
5.2.4. Simulation as an accelerator for alignment
The three alignment mechanisms described above—policies, visualizations, and the coordination agent—improve through repeated interaction. But in live deployment, edge cases surface slowly: a speaker may take days to reply, and unusual requests arise only once across an entire series. Simulation accelerates this by compressing weeks of coordination into minutes while maintaining realistic diversity of cases. Participants found simulations realistic (5.90/7 to 6.75/7), useful for surfacing potential problems (5.80/7 to 6.50/7), and confidence in DoubleAgents robustness increased through simulation (5.80/7 to 6.38/7). P3 described: “The simulations felt realistic and true to the person [they] were simulating… it elicited possible problems that would happen in real life… Like people asking for funding or needing to reschedule—it forced me to think about how I would handle that and to create relevant policies. I can trust the system to handle that in the future since it already happens in the simulation.
Simulation served all three mechanisms as a testbed: it provided a safe environment to probe whether agent behavior matched user intent (policies), populated the dashboard with realistic state for practicing oversight (visualizations), and surfaced ambiguous situations where the agent should pause (stop hooks)—letting participants define delegation boundaries in advance rather than discovering them through live mistakes. This parallels unit testing in software engineering, but where correctness is user-specific rather than universally defined. Simulation is not required—alignment can develop through live use—but it compresses the iteration cycle and lets participants build confidence before real social stakes are involved.